OpenStudy (anonymous):

The measure of the supplement of an angle is 4 times the measure of the angle. What is the measure of each angle?

OpenStudy (anonymous):

original angle = a supplement angle = b 4a = b a + b = 180 you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so substituting 4a for b gives you: a + 4a = 180 can you finish it?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

5a = 180 a = 36 that's your original angle 36 x 4 = 144 that's your supplimentary angle to check your work you can do 36 + 144 which does equal 180 You get it?

OpenStudy (anonymous):

how did you get 36?????

OpenStudy (anonymous):

a + 4a = 5a 5a = 180 divide both sides by 5 a = 36

OpenStudy (anonymous):

when doing problems like this it helps to think of each angle in relation to the other angle. you can draw a picture to help you as well.

OpenStudy (anonymous):

I get it now thank you!

OpenStudy (anonymous):

Anytime, good luck with your studies :)